$10000 was compounded annually at 12.5% p.a. and amounted to $52015.80. How many years did it take?


1
Expert's answer
2021-11-15T16:11:45-0500

"A=P(1+r)^t"


"52015.80=10000(1+0.125)^t"


"5.201580=[1+0.125]^t"


"5.201580=[1.125]^t"


"log5.201580=tlog(1.125)"


"t=\frac{log5.201580}{log1.125}=14years"